OpenStudy (anonymous):

what is the answer to x-1=-5 ?

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

okay thanks , what about -89=n-20

OpenStudy (anonymous):

-69

OpenStudy (anonymous):

x+7=24 ?

OpenStudy (anonymous):

17

OpenStudy (anonymous):

you got any more

OpenStudy (anonymous):

yea , -b/3=12

OpenStudy (anonymous):

36

OpenStudy (anonymous):

or 4x=-236 ?

OpenStudy (anonymous):

-59

OpenStudy (anonymous):

could i have the medals